# Designing for Distributed Heterogeneous Modularity: On Software Architecture and Deployment of the MoonBots

### v1

@@ I Introduction
Achieving true modularity in robotics would mark a significant leap forward [1, 2], especially in the context of space exploration [3, 4, 5]. Most robotic systems today are engineered as monolithic, task-specific platforms, requiring the task or environment to adapt to the robot rather than the opposite [6, 7, 8]. This rigid coupling between form and function significantly limits adaptability.
@@
As a concrete example, no single robot can palletize boxes on an assembly line, drive on public roads, and traverse mountainous terrain. While excellent technologies exist for each task individually [9, 10, 6], they remain siloed and cannot be unified into a single, adaptive system. Humans, in contrast, excel at switching between domains – a capability that continues to inspire the pursuit of general-purpose humanoid robots [11]. But, as illustrated in Fig. 1, modular robots offer an alternative vision: breaking the monolith to build reconfigurable systems capable of adapting to the task, environment, and mission [4, 1, 2].
@@
Our project: MoonShot [12], envisions a fleet of MoonBot robots building lunar infrastructures. As illustrated Figs. 1 and 2 and described in our article [13], MoonBots reconfigure in a multitude of shapes, specialized in diverse tasks – exploration, mapping, construction, and repair – and distributed over wide networks. Modularity is key for such space projects [5, 14]: Redundancy is improved with replaceable parts; weight is reduced by sending modules instead of specialized robots; load is balanced between servers; the system can be upgraded through the years [3, 4].
@@
To proceed, we must clarify our broader notion of modularity. While existing robotic frameworks define it through physical reconfiguration [15, 16], our approach distinguishes itself by not tying modularity to physical shape. In the literature we find three emergent family of modularity: “graft modularity” (augmenting capability by adding components [7, 8]), “clone modularity” (scaling via identical units [2, 17]), and “swarm modularity” (emergent behavior from many similar agents [10, 18]). Our framework aim for a superset, novel to robotics, encompassing all those approaches: “distributed heterogeneous modularity,” where modules differ in form, function, and communication, and are distributed across both physical and computational infrastructure.
@@
We can find our large-scale modularity practiced in space systems [3]. The International Space Station exemplifies this mindset with heterogeneous physical modules, but also distributed components across orbit and ground, cooperating for control, analysis, and fault recovery.
@@
Distributed heterogeneous modularity is now widespread in modern software, with Internet of Tings (IoT) and cloud systems [19]. These domains have developed mature, scalable patterns for managing complexity across heterogeneous layers of infrastructure. We observe those technologies – DDS, MQTT, and Zenoh [20, 21, 22] – arriving in the field of robotics, giving birth to Industrial IoT (IIoT) [19]. Based on those ideas ROS2 [23] is now largely unavoidable in robotics, successfully implementing distributed computation, standardized messaging and packaging. We will see that ROS2 provides most – but not all – of the solutions we need.
@@ II-A Design Philosophy
As highlighted by Conway [24], product and software architecture converges towards the structure of the group of people making it. In robotics, we find that the software is highly correlated to the shape of the robot.
@@
To avoid such shortcomings, we propose a strong component-based software approach [25]. We will avoid encoding relationships between software components, because this would imply relationship between people and robot shape. Instead, we treat the system as assemblies of loosely-coupled components, where behavior emerges from data.
Core to this model is separation of concerns [26, 25]. Components must not be aware of the internal logic – or even existence – of other components. They should operate solely on resources, exposed via a medium of data. Our architecture should be shaped by data, not by the shape of the robot. It is then, the responsibility of higher-order components – solvers, planners, orchestrators, and AI agents – to compose behavior by acting on those data resources.
@@ II-B Network
As in industrial IoT (IIoT) systems, the network is a critical enabler of modularity [19, 22, 17]. Given our goal Fig. 2 and philosophy Sec. II-A, we require a data transport layer offering strong separation of concerns, while supporting long range and unreliable communications.
Traditional Data Delivery Systems (DDS) are efficient and relatively simple to deploy [20]. However, they are strictly peer-to-peer, thus violate separation of concerns by tightly coupling data flow between all participants.
In contrast, MQTT communicates through a central broker [21]. While this decouples participants, it introduces a single point of latency, failure, and bandwidth constraint.
Zenoh [22] offers the best of both: peer-to-peer, brokered, and mesh communication can coexist. Hence, the network topology can adapt to the mission and robot. Independent software routers manage routing and optimization, making network orchestration the sole responsibility of Zenoh.
@@ II-C Launch and Orchestration
As the number of robot modules grows – each potentially composed of multiple onboard computers – the task of launching the correct processes, with appropriate configuration, on the correct machines becomes exponentially more complex, time-consuming, and error-prone. Our orchestrator [25] addresses this by reducing operator intervention down to three actions per robot: update, build, and start. While not fully automated, this already provides critical reliability and speed during deployment.
@@
    Updater and Build: Resolves dependencies, then uses the Doit build system [27] to resolve a graph of offline installation and compilation tasks.
@@ II-D Components
Our model avoids this by isolating robot-specific logic within the well-defined Injection and Override blocks [26, 25] – both of which are chosen by the orchestrator (Sec. II-C). For example, Injection may be a telemetry system, while Override handles a broken joint.
@@
The highly reusable API subcomponent abstracts low-level communication and safety-critical logic behind a consistent interface. As shown in our previous work [28], actuator synchronization can be robustly implemented without requiring detailed system knowledge – allowing safe operations on any actuators.
@@ II-E Implementation and Test Campaigns
The system was tested and improved gradually by intensive test campaigns: three field tests totaling nine weeks at the JAXA sand field [29]; a five-day experiment in DLR’s LUNA lunar simulation facility (Fig. 1a) [30]; and a six-day campaign at the Osaka 2025 Universal Exposition, with daily operations of ten hours. Referring to Fig. 2, these tests implemented MoonBase #1, where multiple students simultaneously operated multiple robots locally. Some tests also involved remote operations, simulating a lunar base controlling robots from afar. In one proof-of-concept trial at the Osaka Exposition, we demonstrated precise teleoperation across multiple wireless links: Operator $\rightarrow$ WiFi $\rightarrow$ Internet $\rightarrow$ Cellular 5G $\rightarrow$ WiFi $\rightarrow$ Robots.
@@ IV-C Applicability
Furthermore, our most fundamental components – joint manager, kinematic manager, API [28], operator interface – are available open-source in our Motion Stack software (https://github.com/2lian/Motion-Stack).
@@ V Conclusion
While not yet space-hardened, our system was deliberately with generality in mind (Sec. IV-C). Our architecture aims to be agnostic to physical implementations, and thus applicable to industrial, terrestrial, and interplanetary deployments alike. Future work will address space-specific constraints, further promising Zenoh tooling, and improving orchestrator via centralized server. This contribution serves as a practical reference for designing scalable, modular robotic systems.

### v2

@@ I Introduction
Achieving true modularity in robotics would mark a significant leap forward [survey_reconf, Yim2009], especially in the context of space exploration [in_space_assembly, space_reconf2021, colab2020]. Most robotic systems today are engineered as monolithic, task-specific platforms, requiring the task or environment to adapt to the robot rather than the opposite [robust_cheeta2018, saber2021, quatro2014]. This rigid coupling between form and function significantly limits adaptability.
@@
As a concrete example, no single robot can palletize boxes on an assembly line, drive on public roads, and traverse mountainous terrain. While excellent technologies exist for each task individually [MoveIt2019, self_driving, robust_cheeta2018], they remain siloed and cannot be unified into a single, adaptive system. Humans, in contrast, excel at switching between domains – a capability that continues to inspire the pursuit of general-purpose humanoid robots [humanoid]. But, as illustrated in Fig. 1, modular robots offer an alternative vision: breaking the monolith to build reconfigurable systems capable of adapting to the task, environment, and mission [space_reconf2021, survey_reconf, Yim2009].
@@
Our project: MoonShot [moonbot_webpage], envisions a fleet of MoonBot robots building lunar infrastructures. As illustrated Figs. 1 and 2 and described in our article [moonbot], MoonBots reconfigure in a multitude of shapes, specialized in diverse tasks – exploration, mapping, construction, and repair – and distributed over wide networks. Modularity is key for such space projects [colab2020, space_busi_2020]: Redundancy is improved with replaceable parts; weight is reduced by sending modules instead of specialized robots; load is balanced between servers; the system can be upgraded through the years [in_space_assembly, space_reconf2021].
@@
To proceed, we must clarify our broader notion of modularity. While existing robotic frameworks define it through physical reconfiguration [moobot_zero, ModularReconfig2020], our approach distinguishes itself by not tying modularity to physical shape. In the literature we find three emergent family of modularity: “graft modularity” (augmenting capability by adding components [saber2021, quatro2014]), “clone modularity” (scaling via identical units [Yim2009, decentra_robot_mod]), and “swarm modularity” (emergent behavior from many similar agents [self_driving, swarm]). Our framework aim for a superset, novel to robotics, encompassing all those approaches: “distributed heterogeneous modularity,” where modules differ in form, function, and communication, and are distributed across both physical and computational infrastructure.
@@
We can find our large-scale modularity practiced in space systems [in_space_assembly]. The International Space Station exemplifies this mindset with heterogeneous physical modules, but also distributed components across orbit and ground, cooperating for control, analysis, and fault recovery.
@@
Distributed heterogeneous modularity is now widespread in modern software, with Internet of Tings (IoT) and cloud systems [iiot]. These domains have developed mature, scalable patterns for managing complexity across heterogeneous layers of infrastructure. We observe those technologies – DDS, MQTT, and Zenoh [DDS, MQTT, zenoh] – arriving in the field of robotics, giving birth to Industrial IoT (IIoT) [iiot]. Based on those ideas ROS2 [ros2] is now largely unavoidable in robotics, successfully implementing distributed computation, standardized messaging and packaging. We will see that ROS2 provides most – but not all – of the solutions we need.
@@ II-A Design Philosophy
As highlighted by Conway [Conway1967HOWDC], product and software architecture converges towards the structure of the group of people making it. In robotics, we find that the software is highly correlated to the shape of the robot.
@@
To avoid such shortcomings, we propose a strong component-based software approach [fundamentals_of_prog]. We will avoid encoding relationships between software components, because this would imply relationship between people and robot shape. Instead, we treat the system as assemblies of loosely-coupled components, where behavior emerges from data.
Core to this model is separation of concerns [programming2007, fundamentals_of_prog]. Components must not be aware of the internal logic – or even existence – of other components. They should operate solely on resources, exposed via a medium of data. Our architecture should be shaped by data, not by the shape of the robot. It is then, the responsibility of higher-order components – solvers, planners, orchestrators, and AI agents – to compose behavior by acting on those data resources.
@@ II-B Network
As in industrial IoT (IIoT) systems, the network is a critical enabler of modularity [iiot, zenoh, decentra_robot_mod]. Given our goal Fig. 2 and philosophy Sec. II-A, we require a data transport layer offering strong separation of concerns, while supporting long range and unreliable communications.
Traditional Data Delivery Systems (DDS) are efficient and relatively simple to deploy [DDS]. However, they are strictly peer-to-peer, thus violate separation of concerns by tightly coupling data flow between all participants.
In contrast, MQTT communicates through a central broker [MQTT]. While this decouples participants, it introduces a single point of latency, failure, and bandwidth constraint.
Zenoh [zenoh] offers the best of both: peer-to-peer, brokered, and mesh communication can coexist. Hence, the network topology can adapt to the mission and robot. Independent software routers manage routing and optimization, making network orchestration the sole responsibility of Zenoh.
@@ II-C Launch and Orchestration
As the number of robot modules grows – each potentially composed of multiple onboard computers – the task of launching the correct processes, with appropriate configuration, on the correct machines becomes exponentially more complex, time-consuming, and error-prone. Our orchestrator [fundamentals_of_prog] addresses this by reducing operator intervention down to three actions per robot: update, build, and start. While not fully automated, this already provides critical reliability and speed during deployment.
@@
    Updater and Build: Resolves dependencies, then uses the Doit build system [doit] to resolve a graph of offline installation and compilation tasks.
@@ II-D Components
Our model avoids this by isolating robot-specific logic within the well-defined Injection and Override blocks [programming2007, fundamentals_of_prog] – both of which are chosen by the orchestrator (Sec. II-C). For example, Injection may be a telemetry system, while Override handles a broken joint.
@@
The highly reusable API subcomponent abstracts low-level communication and safety-critical logic behind a consistent interface. As shown in our previous work [ms_api], actuator synchronization can be robustly implemented without requiring detailed system knowledge – allowing safe operations on any actuators.
@@ II-E Implementation and Test Campaigns
The system was tested and improved gradually by intensive test campaigns: three field tests totaling nine weeks at the JAXA sand field [jaxafield]; a five-day experiment in DLR’s LUNA lunar simulation facility (Fig. 1a) [lunafield]; and a six-day campaign at the Osaka 2025 Universal Exposition, with daily operations of ten hours. Referring to Fig. 2, these tests implemented MoonBase #1, where multiple students simultaneously operated multiple robots locally. Some tests also involved remote operations, simulating a lunar base controlling robots from afar. In one proof-of-concept trial at the Osaka Exposition, we demonstrated precise teleoperation across multiple wireless links: Operator $\rightarrow$ WiFi $\rightarrow$ Internet $\rightarrow$ Cellular 5G $\rightarrow$ WiFi $\rightarrow$ Robots.
@@ IV-C Applicability
Furthermore, our most fundamental components – joint manager, kinematic manager, API [ms_api], operator interface – are available open-source in our Motion Stack software (https://github.com/2lian/Motion-Stack).
@@ V Conclusion
While not yet space-hardened, our system was deliberately designed with generality in mind (Sec. IV-C). Our architecture aims to be agnostic to physical implementations, and thus applicable to industrial, terrestrial, and interplanetary deployments alike. Future work will address space-specific constraints, further promising Zenoh tooling, and improving orchestrator via centralized server. This contribution serves as a practical reference for designing scalable, modular robotic systems.